In 3 h 24 min, a balloon drifts 8.7 km north, 9.7 km east, and 2.9 km in elevation from its release point on the ground. Find (a) the magnitude of its average velocity and (b) the angle its average velocity makes with the horizontal.

Grade:11

1 Answers

Saurabh Kumar
9 years ago
Average velocity is the net displacement covered in the given time.
Using pythagorus theorem, sq rt of (8.7)2 + (9.7)2 = z (let.)
then calculate sq rt (2.9)2 + (x)2 = L
Vavg = L/t